Maintaining constant pressure, the volume of a gas is increased from 18.0 L to 32.0 L by heating it. If the original temperature was 18.0°C, what is the new temperature in degrees Kelvin? Reminder, convert your temperature to Kelvin before you begin the problem.


1
Expert's answer
2022-03-09T07:52:04-0500

P1 = P2 = constant

V1 = 18.0 L

T1 = 18.0°C + 273.15 = 291.15 K

V2 = 32.0 L

T2 = unknown


Solution:

Since the pressure and amount of gas remain unchanged, Charles's law can be used.

Charles's law can be expressed as: V1 / T1 = V2 / T2

V1T2 = V2T1

To find the final temperature, solve the equation for T2:

T2 = V2T/ V1

T2 = (32.0 L × 291.15 K) / (18.0 L) = 517.6 K

T2 = 517.6 K


Answer: The new temperature is 517.6 K